Assignment One

Over the past twenty years, Australian historians have constructed a story of widespread massacres on the frontiers of the expanding pastoral industry that carried off between five and twenty per cent of the Aboriginal population on the mainland … However, (Philipson, 2008) when it is closely examined, the evidence for these claims turns out to be highly suspect. Most of it is poorly founded, other parts are seriously mistaken, and some of it is outright fabrication.

References

Philipson, G. (2008). Imagine a world without copyright. The Sydney Morning Herald, Tuesday June 24, 27.